OpenStudy (anonymous):

What is the best approximation of the projection of (5,-1) onto (2,6)? A. 0.15(2,6) B. 0.10(2,6) C. 0.15(5,-1) D. 0.10(5,-1) @ash2326 :)

OpenStudy (amistre64):

find the cosine of the angle between them, find the length of (5,-1), and unit out the (2,6)

OpenStudy (anonymous):

how do i do that?

OpenStudy (amistre64):

\[proj_u~v\] \[cos(a)=\frac{u\cdot v}{|u||v|}\] \[|u|cos(a)=\frac{u\cdot v}{|v|}\] this is the length of the projection the direction is in the same direction as v, but we have to unit v by dividing out its length \[\frac{|u|cos(a)}{|v|}=\frac{u\cdot v}{|v|^2}\]

OpenStudy (anonymous):

okay :)

OpenStudy (amistre64):

so we already know it has to go in the direction of 2,6 so that narrows it down

OpenStudy (anonymous):

so its between A and B then?

OpenStudy (amistre64):

all you really have to determine now is u.v and |v|

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

how do i determine u x v and IvI ? and the IvI is definitely positive though right? cuz of the absolute value bars?

OpenStudy (amistre64):

|v| = sqrt( vx^2 + vy^2) |v|^2 = vx^2 + vy^2

OpenStudy (amistre64):

square the parts and add em ...

OpenStudy (anonymous):

is vx= 5 and vy=-1?

OpenStudy (amistre64):

no, thats "u" :)

OpenStudy (anonymous):

ah darn :( okay so vx=2 and vy=6 ?

OpenStudy (amistre64):

ye

OpenStudy (anonymous):

so 2^2+6^2= 4+36 = 40 ?

OpenStudy (amistre64):

thats our denominator so far so good

OpenStudy (anonymous):

okay :) yay!

OpenStudy (amistre64):

think of |v|^2 as v.v your just multiplying like component parts and adding u.v is multiplying like component parts and adding

OpenStudy (anonymous):

okay :) i'll try to remember that! so are we looking for the numerator now? :O

OpenStudy (amistre64):

v: 2,6 dot v: 2,6 ----------- 40 = 4+36 u: 5,-1 dot v: 2, 6 ----------- 4 = 10-6

OpenStudy (anonymous):

okay :) so what do i do from there? :O

OpenStudy (anonymous):

|u|cos(a)|v|=u⋅v|v|2 do i apply it to this? the equation u put above^^ ?

OpenStudy (anonymous):

sorry, not good at latex :(

OpenStudy (amistre64):

we just determined all the parts needed: u.v = 4 |v|^2 = 40 and we already know the direction of v is required: what then is 4/40?

OpenStudy (anonymous):

okay.. 4/40= .1 ?

OpenStudy (amistre64):

then all thats left is to pick the correct option

OpenStudy (anonymous):

so the answer is B. 0.10(2,6) ?? :O

OpenStudy (amistre64):

thats what i would go with yes

OpenStudy (anonymous):

awesome!! :) thank you so much!!! :D

OpenStudy (amistre64):

youre welcome, and good luck

OpenStudy (anonymous):

thank you!! :)